# Spin-charge separation and resonant valence bond spin liquid in a frustrated doped Mott insulator

### v1

@@ A.5 Vanishing wave function in the stoichiometric limit
Finally, we study the holon-holon correlations. The dependence on distance between the two holes is shown in Fig. 2 in the main text, for $t=1$ and $J=0$ and system sizes 16 and 32 (and, consistently with the structure of the wave function, it overlaps with the spinon-holon correlation for a single-hole GS). For the 16-site system, we also study the behaviour in presence of interactions, as a function of $J/t$, in Fig. S5. Note that because of the small system size, we only have access to three distances (with periodic boundary conditions, two sites in this system are either one or two lattice spacings apart). Additionally the second-nearest neighbour and third-nearest neighbour correlators give the same values, and we therefore only have two independent values of the correlator. Within the limitation of the small system size accessed in our study, we observe no sign of holon attraction (rather a small tendency to holon repulsion) in the RVB liquid phase. In the large-$J$ ferromagnetic regime, when the GS becomes partially magnetised, holon attraction is introduced by the holons sharing minority spins [Zhang2018].
SM Note F Experiments
@@ F.1 Measurements
As discussed earlier, some studies of this kind have already been performed, including measurements of polycrystalline samples of Y2-xCaxIr2O7 for $x=0$–0.2  [Fukazawa2002, Zhu2014]. Using X-ray photoelectron spectroscopy, Ref. Zhu2014 finds that the amount of Ir5+ in the sample does indeed increase with Ca doping and, at the same time, transport measurements indicate that the conductivity improves. Moreover, a feature in the resistivity that might be associated with the metal-insulator transition seen in other pyrochlore iridates moves to lower temperatures. And temperature-dependent magnetic susceptibility measurements show that the bifurcation of the zero-field and field-cooled traces (which likely arises due to weak ferromagnetic behaviour and is typically associated with the onset of AIAO order in the pyrochlore iridates [Ishikawa2012]), shifts up in temperature with doping. This last result, however, is in contradiction to Ref. Fukazawa2002, which finds that the bifurcation moves to lower temperatures over the same doping range. That study also finds an enhancement of the conductivity as $x$ increases, but no features related to any metal-insulator transition could be discerned. One drawback of these studies is the polycrystalline nature of the samples. Broken symmetry at the surface and the unknown effects of intergrain boundaries in pressed pellets are known to affect the experimental results in these materials, particularly transport measurements, and tend to obscure what is happening in the bulk [Coak2024]. So, while a marked effect on the magnetic and transport properties of pyrochlore iridates has already been observed on doping with holes, measurements of high-quality single crystals are necessary to understand these changes and test the predictions discussed here. This requires further investigation into improving synthesis techniques.
F.2 Crystal growth

### v2

@@ A.5 Vanishing wave function in the stoichiometric limit
Finally, we study the holon-holon correlations. The dependence on distance between the two holes is shown in Fig. 2 in the main text, for $t=1$ and $J=0$ and system sizes 16 and 32 (and, consistently with the structure of the wave function, it overlaps with the spinon-holon correlation for a single-hole GS).
For the 16-site system, we also study the behaviour in presence of interactions, as a function of $J/t$, in Fig. S6. Note that because of the small system size, we only have access to three distances (with periodic boundary conditions, two sites in this system are either one or two lattice spacings apart). Additionally the second-nearest neighbour and third-nearest neighbour correlators give the same values, and we therefore only have two independent values of the correlator. Within the limitation of the small system size accessed in our study, we observe no sign of holon attraction (rather a small tendency to holon repulsion) in the RVB liquid phase. In the large-$J$ ferromagnetic regime, when the GS becomes partially magnetised, holon attraction is introduced by the holons sharing minority spins [Zhang2018].
SM Note F Checkerboard
Figure S7: Holon correlations on the checkerboard lattice. The dark-blue dots show the holon density for a checkerboard lattice of 16, 24, 36 and 48 sites, with $t=1$ and $J=0$ doped with a single hole, as a function of distance from a pinned spin, indicating spinon-holon deconfinement. Crosses show the holon-holon correlations in equivalent systems with two holes; they are numerically identical to the holon-spinon correlations.
In Fig. S7, we show the spinon-holon correlations for checkerboard systems of 16, 24, 36, and 48 sites, with one hole and a pinned spin ($t=1$ and $J=0$). We include in the figure the holon-holon correlations when two holes are present, for the same system sizes. Once again, we find that the holon-holon correlations exactly match the spinon-holon correlations, consistently with our ground state wave function ansatz. Indeed, as already mentioned earlier, we find that the numerically obtained GS wave function for the 16- and 24-site checkerboard lattice is identical within numerical accuracy to a state constructed as the equal amplitude superposition of dimer-singlet states with one or two delocalised holons (with a fixed spin in the one-holon case) and the further constraint of no more than one dimer per tetrahedron.
Figure S8: Spin structure factor for the 36-site checkerboard lattice. This is shown for $t=1$ and $J=0$, doped with two holes. Accessible momenta are marked by grey pluses.
For a checkerboard system of 36-sites and two holes ($t=1$ and $J=0$), we also show in Fig S8 the spin structure factor, for reference.
SM Note G Experiments
@@ G.1 Measurements
As discussed earlier, some studies of this kind have already been performed, including measurements of polycrystalline samples of Y2-xCaxIr2O7 for x=0x=0–0.2  [Fukazawa2002, Zhu2014]. Using X-ray photoelectron spectroscopy, Ref. Zhu2014 finds that the amount of Ir5+ in the sample does indeed increase with Ca doping and, at the same time, transport measurements indicate that the conductivity improves. Moreover, a feature in the resistivity that might be associated with the metal-insulator transition seen in other pyrochlore iridates moves to lower temperatures. And temperature-dependent magnetic susceptibility measurements show that the bifurcation of the zero-field and field-cooled traces (which likely arises due to weak ferromagnetic behaviour and is typically associated with the onset of AIAO order in the pyrochlore iridates [Ishikawa2012]), shifts up in temperature with doping. This last result, however, is in contradiction to Ref. Fukazawa2002, which finds that the bifurcation moves to lower temperatures over the same doping range. That study also finds an enhancement of the conductivity as $x$ increases, but no features related to any metal-insulator transition could be discerned. One drawback of these studies is the polycrystalline nature of the samples. Broken symmetry at the surface and the unknown effects of intergrain boundaries in pressed pellets are known to affect the experimental results in these materials, particularly transport measurements, and tend to obscure what is happening in the bulk [Coak2024]. So, while a marked effect on the magnetic and transport properties of pyrochlore iridates has already been observed on doping with holes, measurements of high-quality single crystals are necessary to understand these changes and test the predictions discussed here. This requires further investigation into improving synthesis techniques.
G.2 Crystal growth